A gas mixture contains O2, N2, and Ar at partial pressures of 125, 175, and 225 mmHg,


respectively. If CO2 gas is added to the mixture until the total pressure reaches 623 mmHg, what is


the partial pressure in milliliters of mercury, of CO2?

1
Expert's answer
2021-11-17T13:37:03-0500

P(O2)=125 mmHg

P(N2)=175 mmHg

P(Ar)=225 mmHg

Then, the total pressure = 623 mm Hg.

Total pressure=P(O2)+P(N2)+P(Ar)+P(CO2)

623 mmHg = (125+175+225+P(CO2)) mmHg

623mmHg=(525+P(CO2)) mmHg

Therefore, P(CO2)=(623−525) mmHg

P(CO2)= 98 mmHg